OpenStudy (anonymous):

A triangle has side lengths 10, 15, and 7. Is the triangle acute, obtuse, or right? Explain

OpenStudy (anonymous):

MY ANSWER: Obtuse, because to find if the triangle is right, acute or obtuse you use the Pythagorean Theorem. You plug in the two smaller numbers in the equation to it looks like this, 7^2+10^2=12.2. Since the hypotenuse given in the problem is larger than the hypotenuse you get in the theorem it is obtuse. Is this right?

OpenStudy (raden):

7, 10, 15 are the sides of triangle. if 7^2 + 10^2 < 15^2 be right, then it is an obtuse. YOU ARE RIGHT

OpenStudy (anonymous):

Really? Okay then i am really confused...Because my teacher said repeatedly that i was wrong?

OpenStudy (raden):

the answer already be right, but your reason not yet :) generally, if given 3 sides of a triangle a, b, and c with a < b < c. so, if a^2 + b^2 < c^2 then it is an obtuse triangle if a^2 + b^2 = c^2 then it is a right triangle if a^2 + b^2 > c^2 then it is a cute triangle

OpenStudy (anonymous):

I see, but she said my work was corrrect but my conclusion wasnt...

OpenStudy (anonymous):

I'm really starting to doubt teachers...

OpenStudy (raden):

yes, because you said it is right triangle. that's the mistake, just check bellow : if given 3 sides of a triangle a, b, and c with a < b < c. so, (1) if a^2 + b^2 < c^2 then it is an obtuse triangle (2) if a^2 + b^2 = c^2 then it is a right triangle (3) if a^2 + b^2 > c^2 then it is a cute triangle

OpenStudy (anonymous):

I said it was obtuse